Steve has some dimes and quarters worth $4.10. He has 8 fewer dimes than quarters, how many of each coins does he have?
Answer by mamiya(56)

let d be the number of dimes and q the number of quarters he has.
So, Steve has $(0.1d + 0.25q)= $4.10
the number of dimes he has is 8 fewer than the number of quarters.
so, d=q-8
from the two equations below, we have the following system:
0.25q + 0.1d= 4.10
q - d = 8

by adding up the first equation and 0.1 time the second equation, we have
0.25q + 0.1d + 0.1q - 0.1d = 4.10 + 0.1*8
0.35q= 4.9
q = 4.9/0.35
= 14
d = q - 8
= 14 - 8
= 6
so, Steve has 14 quarters and 6 dimes.